OpenStudy (anonymous):

What is the product in lowest terms? 1/3 x 1/4 A. 1/12 B. 1/2 C. 1/6 D. 2/7

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

yes it is A.\[\frac{ 1 }{ 3 } * \frac{ 1 }{ 4 }=\frac{ 1*1 }{ 3*4 }=\frac{ 1 }{ 12 }\]